HETAV PATEL

Grade 11,

For a long time an inductor L1 is in series with resistance R in presence of cell of emf 20V. Then at t=0 the inductor L1 is connected in series with inductor L2 and the resistance R with the same cell of emf 20V. Current in the circuit as a function of time is I=2-4/3exp(-2t/3). What is R/L1?

For a long time an inductor L1 is in series with resistance R in presence of cell of emf 20V. Then at t=0 the inductor L1 is connected in series with inductor L2 and the resistance R with the same cell of emf 20V. Current in the circuit as a function of time is I=2-4/3exp(-2t/3). What is R/L1?

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

at t→0,I=20/R,when L2 is connected total flux is conserved,so at t→0, I(L1+L2)=20*L1/R,(dI/dt)(L1+L2)+IR=20,solve this differential equation in I and use the value of current at t=0 to obtain the answer.